OpenStudy (anonymous):

how does this: -3 +5√(2k) ------------------- -2k +4√(5k^2) simplify to this: -3 -6√(5) +5√(2k) +10√(10k) ------------------------------------- 38k

OpenStudy (anonymous):

even a clue based on some web site to read would be helpful .. anything

OpenStudy (anonymous):

Rationalizing the denominator means multiplying by \[\frac{4√(5k^2) }{ 4√(5k^2) }\] I'm assuming this will give you the simplified version.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah I just simplified the first problem on paper, it doesn't simplify to the second one any way I can see

OpenStudy (anonymous):

if i multiplied -3 +5√(2k) -2k +4√(5k^2) ------------------- [multiply by] ------------------- -2k +4√(5k^2) -2k +4√(5k^2) then cancel out "-2k +4√(5k^2) ". I end up with the the same equation i started with. -3 +5√(2k) ------------------- -2k +4√(5k^2)

OpenStudy (anonymous):

the equation to simplify was generated by a windows program that creates work sheets in mathematics. it may be an erroneous answer. i would still like to know how to simplify the equation.

OpenStudy (anonymous):

no no you don't multiply by the -2k when rationalizing the denominator. That's already rational. You multiply by the radical number which is 4sqrt(5x^2) over itself only.

OpenStudy (anonymous):

thanks for attempting hoa :)

OpenStudy (anonymous):

\[\frac{ -3 +5√(2k) }{ -2k +4√(5k^2)} * \frac{ 4√(5k^2)}{ 4√(5k^2) }\] This will rationalize the denominator, which is the only simplification i would make.

OpenStudy (anonymous):

then the resulting multiplication becomes: -3 +5√(2k) 4√(5k^2) ------------------- [multiply by] ------------------- -2k +4√(5k^2) 4√(5k^2) -3 +5√(2k) -------------------------------------- -2k [multiplied by] 4√(5k^2)

OpenStudy (anonymous):

that's wrong. You have to distribute in the numerator, and when you multiply two square roots like you're doing in the denominator it erases the square root symbol.

OpenStudy (anonymous):

Which is the entire point of "rationalizing" usually you don't want a radical number in the denominator.

OpenStudy (anonymous):

i clearly do not know how to do this at all.

OpenStudy (anonymous):

let me work it out on paper and post my own result to rationalizing and you can try and follow the logic

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

this stuff is new to me, i have only been studying for less then a month. i have a test in a couple of weeks.

OpenStudy (anonymous):

\[\frac{ -3+5\sqrt2k }{ -2k+4\sqrt(5k^2) } * \frac{4\sqrt(5k^2) }{ 4\sqrt(5k^2) } =\] \[\frac{ -12\sqrt(5k^2) + 20(\sqrt(10k^3)}{ -2k+16*5k^2} =\] \[\frac{ -12\sqrt(5k^2) + 20(\sqrt(10k^3)}{ -2k+80k^2} =\] Something like this, I'm sure I made a computation error somewhere but this is basically it. You distribute in the numerator and the square root cancels out in the bottom because obviously \[\sqrt(2) * \sqrt(2) = 2\]

OpenStudy (anonymous):

I should have distributed in the bottom as well, I'm retarded. Hopefully you get the principal though.

OpenStudy (anonymous):

yeah, it makes sense :) i have never multiplied square-roots before. i had not known it was possible until i reviewed your example. would the bottom distribute out to be : -8k√(5k^2) + 80k^2